# Enhancing Automatic Chord Recognition via Pseudo-Labeling and Knowledge Distillation

### 26 Mar 2026 v2

@@ Abstract
Automatic Chord Recognition (ACR) is constrained by the scarcity of aligned chord labels, as well-aligned annotations are costly to acquire. At the same time, open-weight pre-trained models are currently more accessible than their proprietary training data. In this work, we present a two-stage training pipeline that leverages pre-trained models together with unlabeled audio. The proposed method decouples training into two stages. In the first stage, we use a pre-trained BTC model [23] as a teacher to generate pseudo-labels for over 1,000 hours of diverse unlabeled audio and train a student model solely on these pseudo-labels. In the second stage, the student is continually trained on ground-truth labels as they become available. To prevent catastrophic forgetting of the representations learned in the first stage, we apply selective knowledge distillation (KD) from the teacher as a regularizer. In our experiments, two models (BTC, 2E1D) were used as students. In stage 1, using only pseudo-labels, the BTC student achieves over 98% of the teacher’s performance, while the 2E1D model achieves about 96% across seven standard mir_eval metrics. After a single training run for both students in stage 2, the resulting BTC student model surpasses the traditional supervised learning baseline by 2.5% and the original pre-trained teacher model by 1.55% on average across all metrics. The resulting 2E1D student model improves over the traditional supervised learning baseline by 2.67% on average and achieves almost the same performance as the teacher. Both cases show large gains on rare chord qualities.
@@ 1 Introduction
We also introduce a compact, dual-encoder architecture (2E1D) that is based on the Transformer architecture [32] and is more lightweight than the teacher model. We demonstrate that our method can generalize across architectures. Our experiments show that the best resulting student surpasses both the traditional supervised learning training method and the pre-trained teacher model across all seven standard mir_eval metrics [26], with particularly large gains on rare chord qualities. A chord recognition web application is open-sourced  [^1] https://github.com/ptnghia-j/ChordMiniApp, available for practitioners to experiment with and validate the performance of our models with their music audio locally.
@@ 3.2 Training Method
We employ a pre-trained teacher model $f_{t}:\mathbb{R}^{T\times F}\rightarrow\mathbb{R}^{T\times C}$ trained on $\mathcal{D}_{l}$ to generate pseudo-labels for unlabeled data. Any open-weight ACR model can serve as the teacher. In our experiments, we use the BTC model of Park et al. [23] as the teacher model, and adopt the standard vocabulary size $C=170$ for all training settings. For each unlabeled sequence $x_{u}\in\mathcal{D}_{u}$, pseudo-labels are generated via frame-wise argmax over teacher outputs:
$\hat{y}_{u}^{(t)}=\arg\max_{c\in\{1,\ldots,C\}}\bigl[f_{t}(x_{u})\bigr]_{t,c},\qquad t=1,\ldots,T$ (1)
@@
Figure 3: Illustration of the proposed two-stage training pipeline. Details of audio data are described in Section 4.1. Note: The resulting Student model CL of stage 2 can be continually trained when additional labeled data is available.
3.2.2 The Pipeline of Training
Figure 3 illustrates the complete two-stage training pipeline. In Stage 1, unlabeled data is first preprocessed into Constant-Q Transform spectrograms following the procedure described in Section 4.1. A pre-trained model is used as a teacher to infer frame-wise pseudo-labels (PL) on the unlabeled audio for each track, producing paired data $(\text{Spectrogram},\text{PL})$. A student model is then trained on these pseudo-labeled pairs until convergence. Optionally, knowledge distillation (KD) from the teacher’s soft targets can be applied during this stage to accelerate convergence; this optional path is depicted as a dashed line in the figure. We denote the resulting model after the first stage of training as the Student model PL. Stage 2 begins when the labeled data becomes available. Labeled audio undergoes the same preprocessing, yielding paired data $(\text{spectrogram},\text{GT})$, where GT denotes ground-truth chord annotations. In this stage, Student model PL serves as the weight initialization for a new model called Student model CL. Then the Student model CL serves as the initialization for continual learning as additional labeled datasets are acquired. The selective KD signal, mentioned in Section 3.2.4, from the original pre-trained teacher is incorporated throughout the training on Student model CL. KD acts as regularization, anchoring the student to the teacher’s distributional knowledge when ground-truth labels conflict with teacher predictions, while still allowing adaptation when they agree. In both stages, the training loss is a weighted sum of a classification term and a KD term as follows:
@@ 3.2.2 The Pipeline of Training
where $\mathcal{D}_{u}^{(p)}$ denotes the pseudo-labeled unlabeled data and $\mathcal{D}_{l}$ denotes the ground-truth labeled data. Setting $\alpha{=}0$ reduces Eq. (3) to pure classification training without KD. The KD loss is defined as:
$\mathcal{L}_{\text{KD}}=\tau^{2}\cdot D_{\text{KL}}\!\left(\sigma\!\left(\frac{\mathbf{z}_{t}}{\tau}\right)\Big\|\sigma\!\left(\frac{\mathbf{z}_{s}}{\tau}\right)\right)$ (5)
@@ 3.2.4 Selective KD
where $\theta_{min}$ filters out uninformative low-confidence samples, $\theta_{max}$ down-weights overconfident predictions that may bias toward majority classes, and factor $K\leq 1$ controls how overconfident predictions are down-weighted. The weighted KD loss becomes $\mathcal{L}_{KD}^{sel}=\frac{1}{N}\sum_{i}w(c_{i})\cdot\ell_{KD}(x_{i})$. Samples near decision boundaries (moderate confidence) contain valuable uncertainty information, so we preserve full weight in the informative range $\theta_{min},\theta_{max}]$. We apply selective KD throughout all Stage 2 continual learning experiments to stabilize training by reducing gradient variance from extreme-confidence samples. We set the hyperparameters with $\theta_{min}=0.1,\theta_{max}=0.9,K=0.8$ as robust defaults based on the observed distribution of the teacher confidence.
@@ 4.2 Training Configuration
For the supervised learning (SL) baseline, both BTC and 2E1D are trained from scratch on the full labeled training set (420 songs). The SL baseline uses the same configuration as in pseudo-labeling training, except for the learning schedule. The learning rate decays at a rate of 0.95 when validation accuracy does not improve. Pitch-shifting augmentation via the Rubber Band library [5] transposes both audio and labels from 5-5 to +6+6 semitones.
@@ 4.3 Evaluation Metrics
We employ standard metrics from mir_eval library [26]. The Root metric compares the root note. The Thirds metric adds major and minor third intervals. The Triads metric evaluates all triadic qualities, while Majmin focuses on major and minor qualities. The Sevenths metric measures a predefined set of seventh chords. The Tetrads metric extends evaluation to four tones. MIREX considers an estimation accurate when at least three pitch classes are correct. Frame-wise accuracy (Acc), precision (Prec), and recall (Rec) are computed using standard True Positive, False Positive, and False Negative counts. Additionally, we report Chord Symbol Recall (CSR) and Weighted Chord Symbol Recall (WCSR) following [1]:
@@
where for a chord class $c$ and track index $i$, $S_{c,i}^{pred}$ denotes predicted chord labels, $S_{c,i}^{ref}$ represents ground-truth labels, and $T_{i}$ is the duration of the $i$-th track. The Average Chord Quality Accuracy (ACQA), with the chord set $C$, is calculated as:
@@
$\text{Over}=1-\frac{\sum_{I\in\mathcal{I}_{ref}}\max_{J\in\mathcal{I}_{est}}|I\cap J|}{\sum_{I\in\mathcal{I}_{ref}}|I|}$ (11)
@@
$\text{Under}=1-\frac{\sum_{J\in\mathcal{I}_{est}}\max_{I\in\mathcal{I}_{ref}}|I\cap J|}{\sum_{J\in\mathcal{I}_{est}}|J|}$ (12)
@@
Table 1: Pseudo-labeling results across dataset configurations. The mir_eval metrics are computed against ground-truth test set, while frame-wise accuracy, precision, recall, and F1 measure agreement with the teacher’s predictions. FMA contains short-form clips ($\sim$30s), while DALI and MAESTRO provide long-form full tracks. The pre-trained BTC teacher [23] serves as the upper bound. Best results for BTC and 2E1D student models in the table are highlighted in blue and green, respectively.
@@ 5.1 Stage 1: Training with pseudo-labels
Increasing unlabeled data diversity consistently improves performance across the mir_eval metric hierarchy and segmentation quality (Table 1). Training on all three datasets (FMA, MAESTRO, DALI) yields the strongest results for both architectures, indicating that pseudo-label pretraining benefits from coverage of varied genres and recording conditions. Long-form datasets consisting of full-length tracks (DALI, MAESTRO) produce more stable pseudo-labels than short-form clips (FMA): DALI achieves higher frame-wise agreement with the teacher than FMA. Longer musical contexts provide consistent harmonic progressions, reducing boundary jitter and spurious chord predictions. However, FMA achieves better mir_eval results than DALI for 2E1D due to its larger size. The best BTC student reaches over 98% of teacher performance on ground-truth evaluation across all mir_eval metrics, while the purely transformer-based 2E1D achieves 96–97% of the teacher model results, demonstrating cross-architecture knowledge transfer. When all training targets are pseudo-labels, adding KD better aligns the student to the teacher’s soft-label distribution, resulting in improvements across frame-wise metrics despite a decrease in mir_eval metrics. In addition, KD accelerates optimization, as all pseudo-label training runs using KD converge in 10 epochs versus 20–40 for other settings; this comes with a modest training-time memory overhead to store and backpropagate full soft targets.
@@ 5.2 Ablation: KD as Regularization
To understand KD’s role as a stabilizer under label noise, we evaluate continual learning with misaligned labels at varying $\alpha$ values (Table 2). Without KD ($\alpha{=}0$), both architectures suffer severe and consistent degradation across all metrics: BTC drops substantially throughout the recognition hierarchy, while the wider 2E1D collapses far more drastically to near-failure levels. The dual attention architecture is more susceptible to noisy label overfitting than the deeper CNN-transformer BTC. KD acts as an anchor, preventing noisy fine-tuning from destroying a good pseudo-label initialization. KD adds robustness without impeding adaptation. BTC peaks at $\alpha{=}0.3$ with a broad, consistent recovery across all metrics relative to the no-KD baseline, while 2E1D requires stronger regularization ($\alpha{=}0.5$) to stabilize. Figure 5 visualizes training dynamics. Without KD ($\alpha{=}0$), validation loss rises as the model overfits to noisy labels, whereas when $\alpha>0$}, KD anchors the student to the teacher’s distribution. This supports KD as an effective safeguard that preserves pseudo-label knowledge during adaptation.
@@
mir_eval Metrics Subset of per-Chord Quality Accuracy Aggregate
@@ 5.3 Stage 2: Continually training with GT + KD
Comparison to Prior Work. Table 3 demonstrates that pseudo-label quality is the critical determinant of semi-supervised learning performance. Prior work [1, 16] relies on supervised models trained on limited labeled data to generate pseudo-labels. This drastically constrains label quality and results in weaker model performance. In contrast, training with high-quality pseudo-labels generated by a better teacher model provides better-initialized weights, thereby creating a stronger starting point for model training and adaptation.
@@ 6 Conclusion
Since model weights are often more readily available than proprietary training data, we present a practical training strategy for the ACR problem that leverages open-weight pre-trained models when high-quality labels are scarce. We show that students trained solely on pseudo-labels can approach teacher-level performance across seven mir_eval metrics. We further demonstrate that continual learning can improve performance without catastrophic forgetting when the teacher provides sufficiently general representations. Under our pipeline, the student ultimately surpasses the teacher, especially on rare chord qualities (e.g., Dim, Dim7, Aug). Knowledge distillation improves robustness to noisy labels while preserving adaptability when ground-truth annotations are clean. We also find that the wider 2E1D architecture requires stronger KD regularization than the deeper BTC, underscoring how architectural choices influence continual-learning stability. A key limitation is reliance on teacher quality: biased or weakly generalizable teacher representations can transfer these shortcomings to the student. Future work will explore stronger teacher models and model architectures, ensembles of multiple teachers, scaling to additional unlabeled corpora, and extending the framework to related MIR tasks such as beat tracking and key estimation.

### 28 Mar 2026 v3

@@ Abstract
Automatic Chord Recognition (ACR) is constrained by the scarcity of aligned chord labels, as well-aligned annotations are costly to acquire. At the same time, open-weight pre-trained models are currently more accessible than their proprietary training data. In this work, we present a two-stage training pipeline that leverages pre-trained models together with unlabeled audio. The proposed method decouples training into two stages. In the first stage, we use a pre-trained BTC model [23] as a teacher to generate pseudo-labels for over 1,000 hours of diverse unlabeled audio and train a student model solely on these pseudo-labels. In the second stage, the student is continually trained on ground-truth labels as they become available. To prevent catastrophic forgetting of the representations learned in the first stage, we apply selective knowledge distillation (KD) from the teacher as a regularizer. In our experiments, two models (BTC, 2E1D) were used as students. In stage 1, using only pseudo-labels, the BTC student achieves over 99% of the teacher’s performance, while the 2E1D model achieves about 97% across seven standard mir_eval metrics. After a single training run for both students in stage 2, the resulting BTC student model surpasses the traditional supervised learning baseline by 2.5% and the original pre-trained teacher model by 1.1-3.2% across all metrics. The resulting 2E1D student model improves over the traditional supervised learning baseline by 2.67% on average and achieves almost the same performance as the teacher. Both cases show large gains on rare chord qualities.
@@ 1 Introduction
We also introduce a compact, dual-encoder architecture (2E1D) that is based on the Transformer architecture [32] and is lighter-weight than the teacher model. We demonstrate that our method can generalize across architectures. Our experiments show that the best resulting student surpasses both the traditional supervised learning training method and the pre-trained teacher model across all seven standard mir_eval metrics [26], with particularly large gains on rare chord qualities. A chord recognition web application is open-sourced  [^1] https://github.com/ptnghia-j/ChordMiniApp, available for practitioners to experiment with and validate the performance of our models with their music audio locally.
@@ 3.2 Training Method
We employ a pre-trained teacher model $f_{t}:\mathbb{R}^{T\times F}\rightarrow\mathbb{R}^{T\times C}$ to generate pseudo-labels for unlabeled data. Any open-weight ACR model can serve as the teacher. In our experiments, we use the BTC model of Park et al. [23] as the teacher model, and adopt the standard vocabulary size $C=170$ for all training settings. For each unlabeled sequence $x_{j}\in\mathcal{D}_{u}$, pseudo-labels are generated via frame-wise argmax over teacher outputs:
$\hat{y}_{j}^{(t)}=\arg\max_{c\in\{1,\ldots,C\}}\bigl[f_{t}(x_{j})\bigr]_{t,c},\qquad t=1,\ldots,T$ (1)
@@
Figure 3: Illustration of the proposed two-stage training pipeline. Details of audio data are described in Section 4.1. Note: The resulting Student model CL from stage 2 can be continually trained when additional labeled data is available.
3.2.2 The Training Pipeline
Figure 3 illustrates the complete two-stage training pipeline. In Stage 1, unlabeled data is first preprocessed into Constant-Q Transform spectrograms following the procedure described in Section 4.1. A pre-trained model is used as a teacher to infer frame-wise pseudo-labels (PL) on the unlabeled audio for each track, producing paired data $(\text{Spectrogram},\text{PL})$. A student model is then trained on these pseudo-labeled pairs until convergence. Optionally, knowledge distillation (KD) from the teacher’s soft targets can be applied during this stage to accelerate convergence; this optional path is depicted as a dashed line in the figure. We denote the resulting model after the first stage of training as the Student model PL. Stage 2 begins when the labeled data becomes available. Labeled audio undergoes the same preprocessing, yielding paired data $(\text{spectrogram},\text{GT})$, where GT denotes ground-truth chord annotations. In this stage, Student model PL serves as the weight initialization for a new model called Student model CL. Then the Student model CL serves as the initialization for continual learning as additional labeled datasets are acquired. The selective KD signal, mentioned in Section 3.2.4, from the original pre-trained teacher is incorporated throughout the training on Student model CL. KD acts as regularization, anchoring the student to the teacher’s distributional knowledge when ground-truth labels conflict with teacher predictions, while still allowing adaptation when they agree. When KD is applied, the training loss is a weighted sum of a classification term and a KD term as follows:
@@ 3.2.2 The Training Pipeline
where $\mathcal{D}_{u}^{(p)}$ denotes the pseudo-labeled unlabeled data and $\mathcal{D}_{l}$ denotes the ground-truth labeled data. Setting $\alpha{=}0$ reduces Eq. (3) to pure classification training without KD. The per-frame KD loss is:
$\ell_{\text{KD}}=\tau^{2}\cdot D_{\text{KL}}\!\left(\sigma\!\left(\frac{\mathbf{z}_{t}}{\tau}\right)\Big\|\sigma\!\left(\frac{\mathbf{z}_{s}}{\tau}\right)\right)$ (5)
@@ 3.2.4 Selective KD
where $\theta_{min}$ filters out uninformative low-confidence samples, $\theta_{max}$ down-weights overconfident predictions that may bias toward majority classes, and factor $K\leq 1$ controls how overconfident predictions are down-weighted. The weighted KD loss becomes $\mathcal{L}_{KD}^{sel}=\frac{1}{N}\sum_{i}w(c_{i})\cdot\ell_{KD}(x_{i})$ over all $N$ frames. Samples near decision boundaries (moderate confidence) contain valuable uncertainty information, so we preserve full weight in the informative range $\theta_{min},\theta_{max}]$. We apply selective KD throughout all Stage 2 continual learning experiments to stabilize training by reducing gradient variance from extreme-confidence samples. We set the hyperparameters with $\theta_{min}=0.1,\theta_{max}=0.9,K=0.8$ as robust defaults based on the observed teacher-confidence distribution.
@@ 4.2 Training Configuration
For the supervised learning (SL) baseline, both BTC and 2E1D are trained from scratch on the full labeled training set (420 songs). The SL baseline uses the same configuration as in pseudo-labeling training, except for the learning schedule. The learning rate decays at a rate of 0.9 when validation accuracy does not improve. Pitch-shifting augmentation via the Rubber Band library [5] transposes both audio and labels from 5-5 to +6+6 semitones.
@@ 4.3 Evaluation Metrics
We employ standard metrics from mir_eval library [26]. The Root metric compares the root note. The Thirds metric adds major and minor third intervals. The Triads metric evaluates all triadic qualities, while Majmin focuses on major and minor qualities. The Sevenths metric measures a predefined set of seventh chords. The Tetrads metric extends evaluation to four tones. MIREX considers an estimation accurate when at least three pitch classes are correct. Frame-wise accuracy (Acc), precision (Prec), recall (Rec), and F1 are computed using standard True Positive, False Positive, and False Negative counts. Additionally, we report Chord Symbol Recall (CSR) and Weighted Chord Symbol Recall (WCSR) following [1]:
@@
where, for a chord class $c$ and track index $i$, $S_{c,i}^{pred}$ denotes predicted chord labels, $S_{c,i}^{ref}$ denotes ground-truth labels for respective intervals, and $T_{i}$ is the duration of the $i$-th track. The Average Chord Quality Accuracy (ACQA), with the chord set $C$, is calculated as:
@@
$\text{Over}=\frac{\sum_{I\in\mathcal{I}_{ref}}\max_{J\in\mathcal{I}_{est}}|I\cap J|}{\sum_{I\in\mathcal{I}_{ref}}|I|}$ (11)
@@
$\text{Under}=\frac{\sum_{J\in\mathcal{I}_{est}}\max_{I\in\mathcal{I}_{ref}}|I\cap J|}{\sum_{J\in\mathcal{I}_{est}}|J|}$ (12)
@@
Table 1: Pseudo-labeling results across dataset configurations. The mir_eval metrics are computed against the ground-truth test set, while frame-wise accuracy, precision, recall, and F1 measure agreement with the teacher’s predictions. FMA contains short-form clips ($\sim$30s), while DALI and MAESTRO provide long-form full tracks. The pre-trained BTC teacher [23] serves as the upper bound. Best results for BTC and 2E1D student models in the table are highlighted in blue and green, respectively.
@@ 5.1 Stage 1: Training with pseudo-labels
Increasing unlabeled data diversity consistently improves performance across the mir_eval metric hierarchy and segmentation quality (Table 1). Training on all three datasets (FMA, MAESTRO, DALI) yields the strongest results for both architectures, indicating that pseudo-label pretraining benefits from coverage of varied genres and recording conditions. Long-form datasets consisting of full-length tracks (DALI, MAESTRO) produce more stable pseudo-labels than short-form clips (FMA): DALI achieves higher frame-wise agreement with the teacher than FMA. Longer musical contexts provide consistent harmonic progressions, reducing boundary jitter and spurious chord predictions. However, FMA achieves better mir_eval results than DALI for 2E1D due to its larger size. The best BTC student reaches over 99% of teacher performance on ground-truth evaluation across all mir_eval metrics, while the purely transformer-based 2E1D achieves 96–98% of the teacher model results, demonstrating cross-architecture knowledge transfer. When all training targets are pseudo-labels, adding KD better aligns the student to the teacher’s soft-label distribution, resulting in improvements across frame-wise metrics despite a decrease in mir_eval metrics. In addition, KD accelerates optimization, as all pseudo-label training runs using KD converge in 30-40 epochs versus 50–70 for other settings; this comes with a modest training-time memory overhead to store and backpropagate full soft targets.
@@ 5.2 Ablation: KD as Regularization
To understand KD’s role as a stabilizer under label noise, we evaluate continual learning with misaligned labels at varying $\alpha$ values (Table 2). Without KD ($\alpha{=}0$), both architectures suffer severe and consistent degradation across all metrics: BTC drops substantially throughout the recognition hierarchy, while the wider 2E1D collapses far more drastically to near-failure levels. The dual attention architecture is more susceptible to noisy label overfitting than the deeper BTC architecture. KD acts as an anchor, preventing noisy fine-tuning from destroying a good pseudo-label initialization. KD adds robustness without impeding adaptation. BTC peaks at $\alpha{=}0.3$ with a broad, consistent recovery across all metrics relative to the no-KD baseline, while 2E1D requires stronger regularization ($\alpha{=}0.5$) to stabilize. Figure 5 visualizes training dynamics. Without KD ($\alpha{=}0$), validation loss rises as the model overfits to noisy labels, whereas when $\alpha>0$}, KD anchors the student to the teacher’s distribution. This supports KD as an effective safeguard that preserves pseudo-label knowledge during adaptation.
@@
mir_eval Metrics Subset of Chord Quality Accuracy Aggregate
@@ 5.3 Stage 2: Continually training with GT + KD
Comparison to Prior Work. Table 3 demonstrates that the quality of pseudo-labels is a key factor in the performance of semi-supervised learning. Prior work [1, 16] relies on supervised models trained on limited labeled data to generate pseudo-labels. This constrains label quality and results in weaker model performance. In contrast, training with high-quality pseudo-labels generated by a better teacher model provides better-initialized weights, thereby creating a stronger starting point for model training and adaptation.
@@ 6 Conclusion
Since model weights are often more readily available than proprietary training data, we present a practical training strategy for the ACR problem that leverages open-weight pre-trained models when high-quality labels are scarce. We show that students trained solely on pseudo-labels can approach teacher-level performance across seven mir_eval metrics. We further demonstrate that continual learning can improve performance without catastrophic forgetting when the teacher provides sufficiently general representations. Under our pipeline, the best student model ultimately surpasses the teacher, especially on rare chord qualities (e.g., Dim, Dim7, Aug). Knowledge distillation improves robustness to noisy labels while preserving adaptability when ground-truth annotations are clean. We also find that the wider 2E1D architecture requires stronger KD regularization than the deeper BTC, underscoring how architectural choices influence continual-learning stability. A key limitation is reliance on teacher quality: biased or weakly generalizable teacher representations can transfer these shortcomings to the student. Future work will explore stronger teacher models and model architectures, ensembles of multiple teachers, scaling to additional unlabeled corpora, and extending the framework to related MIR tasks such as beat tracking and key estimation.